OpenStudy (anonymous):

-3y=2x-7 three points then graph

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

graph?

OpenStudy (anonymous):

yeah i need 3 points then i have to graph it

OpenStudy (anonymous):

Probably you could take the points as x=0,x=1, and x=-1. Find the y values for each particular x value, then plot the graph.

OpenStudy (anonymous):

-3y=2x-7 solve for y y=-2/3x+7/3  down 2 over 3 y=mx+b where b is 7/3 the y intercep m is slope down 2 over 3

OpenStudy (anonymous):

but i need three points to fill in x,y

OpenStudy (anonymous):

its linear equations

OpenStudy (anonymous):

not really. THen pic 3 points lol

OpenStudy (anonymous):